# Linear time-and-space-invariant relaxation systems

### v1

@@ V Impedance passivity and relaxation
In this section, we will show that LTSI systems that are internally of relaxation type are impedance passive with energy storage that can be determined from their input-output behaviour, thus extending the analogous result on LTI relaxation systems [6]. In doing so, we will also characterize impedance passivity for LTSI systems, namely, we will show that it can be verified pointwise, just like (internal) relaxation.
To begin with, impedance passivity is the analogue of passivity for infinite-dimensional systems, see [15, Section 7.5] for details. It is defined as follows.
@@
The LTSI system $\Sigma$ is impedance passive if and only if there exist $\hat{Q}_{\omega}\in\mathbb{C}^{n\times n}$, $\hat{Q}_{\omega}\geq 0$, $\omega\in\mathbb{R}^{s}$, such that
@@
This proof is similar to the proof of [16, Theorem 2]. The idea is that the condition $\sup_{\omega\in\mathbb{R}^{s}}\lVert\hat{Q}_{\omega}\rVert<\infty$ is equivalent to the boundedness of the multiplication operator with symbol $\hat{Q}$, which we denote by $\hat{\mathcal{Q}}\in\mathscr{B}(\mathcal{Z})$. We start by proving sufficiency. Since $\hat{Q}_{\omega}\geq 0$ and (21) hold for all $\omega\in\mathbb{R}^{s}$, it follows that $\hat{\mathcal{Q}}\geq 0$ and
@@
for all $\hat{z}\in\mathcal{D}(\hat{\mathcal{A}})$. Recall that the Fourier transform is unitary. This implies that $\mathcal{Q}=\mathcal{F}^{\ast}\hat{\mathcal{Q}}\mathcal{F}\in\mathscr{B}(\mathcal{Z})$, satisfies $\mathcal{Q}\geq 0$ and the conditions of Lemma 1, hence $\Sigma$ is impedance passive.
@@
$\mathcal{S}(z(0))=\int_{0}^{\infty}\langle u(-t),y(t)\rangle\,\textrm{d}t$ (23)
@@
$\hat{y}(0,\omega)=\int_{-\infty}^{0}\hat{g}(-t,\omega)\hat{u}(t,\omega)\,\textrm{d}t$ (29)
Multiplying the latter by $\hat{y}(t,\omega)^{\ast}$ and rearranging yields
$\lvert\hat{y}(0,\omega)\rvert^{2}=\int_{0}^{\infty}\hat{u}(-t,\omega)\hat{y}(t,\omega)^{\ast}\,\textrm{d}t$ (30)
@@
$\hat{y}(t,\omega)^{\ast}=\hat{g}(t,\omega)\hat{y}(0,\omega)^{\ast}$ (31)
Integrating both sides of (30) yields (28), hence the energy storage of the diffusion equation can be determined the past input, which confirms the result of Theorem 4.
@@ VI The Hankel operator and relaxation
In this subsection, we will show that relaxation can be characterized via the Hankel operator. In particular, we will show that an exponentially stable LTSI system is of relaxation type if and only if its Hankel operator is self-adjoint and nonnegative. Consequently, the Hankel operator of an LTSI relaxation system is cyclically monotone and, thus, the gradient of a closed convex functional, which coincides with the (impedance) passive storage of Theorem 4.
@@
$(\mathcal{H}v)(t)=\int_{\mathbb{R}_{+}}\mathcal{G}(t+\tau)v(\tau)\,\textrm{d}\tau$ (32)
where $\mathscr{D}(\mathcal{H})\subset L_{2}(\mathbb{R}_{+},\mathcal{U})$. The Hankel operator maps past inputs to future outputs in the sense that the input given by $u(t)=v(-t)$ for $t\leq 0$ and $u(t)=0$ for all $t>0$, results in the output given by $y(t)=(\mathcal{H}v)(t)$ for all $t\geq 0$. The domain of $\mathcal{H}$ is nontrivial only if the system is stable. Here, we will assume that $\Sigma$ is exponentially stable, i.e., there exists constants $M,\alpha>0$ such that $\lVert\exp(\mathcal{A}t)\rVert\leq Me^{-\alpha t}$ for all $t\in\mathbb{R}_{+}$, see [15, Section 4.1] for details. Then, we obtain the following theorem.
@@
The operators $\mathcal{B}_{\infty}$ and $\mathcal{C}_{\infty}$ are known as the extended controllability and observability maps, respectively, see [15, Section 6.4] for details. They are well-defined and bounded if $\Sigma$ is exponentially stable, hence $\mathcal{H}=\mathcal{C}_{\infty}\mathcal{B}_{\infty}$ is a bounded linear operator from $L_{2}(\mathbb{R}_{+},\mathcal{U})$ to $L_{2}(\mathbb{R}_{+},\mathcal{Y})$. ∎
In [3], it is shown that a stable LTI system is of relaxation type if and only if its Hankel operator is self-adjoint and nonnegative (hence cyclically monotone). Here, we extend this result to exponentially stable LTSI systems.
@@
Note that $\mathcal{H}\in\mathscr{B}(L_{2}(\mathbb{R}_{+},\mathcal{U}))$ due to Theorem 5 and the assumption that $\Sigma$ is exponentially stable. We first prove necessity. Suppose that $\Sigma$ is of relaxation type. It follows that $\hat{\Sigma}_{\omega}$ is of relaxation type for all $\omega\in\mathbb{R}^{s}$ due to Theorem 1. Furthermore, due to [16, Thoerem 1], $\hat{\Sigma}_{\omega}$ is stable for all $\omega\in\mathbb{R}^{s}$. Consequently, due to [3, Theorem 5], the Hankel operator ^$\hat{\mathcal{H}}_{\omega}\in\mathscr{B}(L_{2}(\mathbb{R}_{+},\mathbb{C}^{m}))$ of $\hat{\Sigma}_{\omega}$ satisfies ^$\hat{\mathcal{H}}_{\omega}\geq 0$ for all $\omega\in\mathbb{R}^{s}$. Note that
^$\langle\hat{\mathcal{H}}_{\omega}\hat{v}_{\omega},\hat{w}_{\omega}\rangle=\int_{\mathbb{R}_{+}}\int_{\mathbb{R}_{+}}\langle\hat{g}_{\omega}(t+\tau)\hat{v}_{\omega}(\tau),\hat{w}_{\omega}(t)\rangle\,\textrm{d}\tau\textrm{d}t$ (36)
@@
where $\hat{v}_{\omega},\hat{w}_{\omega}\in L_{2}(\mathbb{R}_{+},\mathbb{C}^{m})$ for almost all $\omega\in\mathbb{R}^{s}$. Therefore, $\mathcal{H}\geq 0$ because ^$\hat{\mathcal{H}}_{\omega}\geq 0$ for all $\omega\in\mathbb{R}^{s}$.
@@
due to (38) and the assumption that $\mathcal{H}$ is self-adjoint. The latter holds for all $\bar{v},\bar{w}\in L_{2}(\mathbb{R}_{+},\mathbb{C}^{m})$ and compact $\Omega\subset\mathbb{R}^{s}$, hence ^$\hat{\mathcal{H}}_{\omega}$ is self-adjoint for almost all $\omega\in\mathbb{R}^{s}$. Similarly, ^$\hat{\mathcal{H}}_{\omega}$ is nonnegative for almost all $\omega\in\mathbb{R}^{s}$. Due to [3, Theorem 5], this implies that $\hat{\Sigma}_{\omega}$ is of relaxation type and, thus, $\hat{g}_{\omega}$ is completely monotone for almost all $\omega\in\mathbb{R}^{s}$. Note that $\hat{g}_{\omega}(t)$ and its time derivatives are continuous in $\omega\in\mathbb{R}^{s}$ for each $t\in\mathbb{R}_{+}$ because the symbols $\hat{A}$, $\hat{B}$ and $\hat{C}$ are assumed to be continuous. Therefore, by continuity, $\hat{g}_{\omega}$ is completely monotone for all $\omega\in\mathbb{R}^{s}$, hence $\Sigma$ is of relaxation type due to Theorem 1. ∎

### v2

@@ V Impedance passivity and relaxation
In this section, we show that LTSI systems that are internally of relaxation type are impedance passive with energy storage that can be determined from their input-output behaviour, thus extending the analogous result on LTI relaxation systems [6]. In doing so, we also characterize impedance passivity for LTSI systems, namely, we show that it can be verified pointwise, just like (internal) relaxation.
To begin with, impedance passivity is the analogue of passivity for infinite-dimensional systems, see [16, Section 7.5] for details. It is defined as follows.
@@
The LTSI system $\Sigma$ is impedance passive if and only if there exist $\hat{Q}_{\omega}\in\mathbb{C}^{n\times n}$, $\hat{Q}_{\omega}\geq 0$, such that
@@
This proof is similar to the proof of [17, Theorem 2]. The idea is that the condition $\sup_{\omega\in\mathbb{R}^{s}}\lVert\hat{Q}_{\omega}\rVert<\infty$ is equivalent to the boundedness of the multiplication operator with symbol $\hat{Q}$, which we denote by $\hat{\mathcal{Q}}\in\mathscr{B}(\mathcal{Z})$. We start by proving sufficiency. Since $\hat{Q}_{\omega}\geq 0$ and (21) hold for all $\omega\in\mathbb{R}^{s}$, it follows that $\hat{\mathcal{Q}}\geq 0$ and
@@
[…]
@@
$\mathcal{S}(z(0))=\int_{0}^{\infty}\langle u(-t),y(t)\rangle\,\textrm{d}t,$ (23)
@@
$\hat{y}(0,\omega)=\int_{-\infty}^{0}\hat{g}(-t,\omega)\hat{u}(t,\omega)\,\textrm{d}t.$ (29)
Multiplying the latter by $\hat{y}(0,\omega)^{\ast}$ and rearranging yields
$\lvert\hat{y}(0,\omega)\rvert^{2}=\int_{0}^{\infty}\hat{u}(-t,\omega)\hat{y}(t,\omega)^{\ast}\,\textrm{d}t,$ (30)
@@
$\hat{y}(t,\omega)^{\ast}=\hat{g}(t,\omega)\hat{y}(0,\omega)^{\ast}.$ (31)
Integrating both sides of (30) yields (28), hence the energy storage of the diffusion equation can be determined from the past input, which confirms the result of Theorem 4.
@@ VI The Hankel operator and relaxation
In this subsection, we show that relaxation can be characterized via the Hankel operator. In particular, we show that an exponentially stable LTSI system is of relaxation type if and only if its Hankel operator is self-adjoint and nonnegative. Consequently, the Hankel operator of an LTSI relaxation system is cyclically monotone and, thus, the gradient of a closed convex functional, which coincides with the (impedance) passive storage of Theorem 4.
@@
$(\mathcal{H}v)(t)=\int_{\mathbb{R}_{+}}\mathcal{G}(t+\tau)v(\tau)\,\textrm{d}\tau,$ (32)
where $\mathscr{D}(\mathcal{H})\subset L_{2}(\mathbb{R}_{+},\mathcal{U})$. The Hankel operator maps past inputs to future outputs in the sense that the input given by u(t)=v(t)u(t)=v(-t) for $t\leq 0$ and $u(t)=0$ for all $t>0$, results in the output given by $y(t)=(\mathcal{H}v)(t)$ for all $t\geq 0$. The domain of $\mathcal{H}$ is nontrivial only if the system is stable. Here, we assume that $\Sigma$ is exponentially stable, i.e., there exists $M,\alpha>0$ such that $\lVert\exp(\mathcal{A}t)\rVert\leq Me^{-\alpha t}$ for all $t\in\mathbb{R}_{+}$, see [16, Section 4.1] for details. Then, we obtain the following theorem.
@@
The operators $\mathcal{B}_{\infty}$ and $\mathcal{C}_{\infty}$ are known as the extended controllability and observability maps, respectively, see [16, Section 6.4] for details. They are well-defined and bounded if $\Sigma$ is exponentially stable, hence $\mathcal{H}=\mathcal{C}_{\infty}\mathcal{B}_{\infty}$ is a bounded linear operator from $L_{2}(\mathbb{R}_{+},\mathcal{U})$ to $L_{2}(\mathbb{R}_{+},\mathcal{Y})$. ∎
As shown in [3], a stable LTI system is of relaxation type if and only if its Hankel operator is self-adjoint and nonnegative (hence cyclically monotone). Here, we extend this result to exponentially stable LTSI systems.
@@
Note that $\mathcal{H}\in\mathscr{B}(L_{2}(\mathbb{R}_{+},\mathcal{U}))$ due to Theorem 5 and the assumption that $\Sigma$ is exponentially stable. We first prove necessity. Suppose that $\Sigma$ is of relaxation type. It follows that $\hat{\Sigma}_{\omega}$ is of relaxation type for all $\omega\in\mathbb{R}^{s}$ due to Theorem 1. Furthermore, due to [17, Theorem 1], $\hat{\Sigma}_{\omega}$ is stable for all $\omega\in\mathbb{R}^{s}$. Consequently, due to [3, Theorem 5], the Hankel operator $\hat{\mathcal{H}}_{\omega}\in\mathscr{B}(L_{2}(\mathbb{R}_{+},\mathbb{C}^{m}))$ of $\hat{\Sigma}_{\omega}$ satisfies $\hat{\mathcal{H}}_{\omega}\geq 0$ for all $\omega\in\mathbb{R}^{s}$. Note that
$\langle\hat{\mathcal{H}}_{\omega}\hat{v}_{\omega},\hat{w}_{\omega}\rangle=\int_{\mathbb{R}_{+}}\int_{\mathbb{R}_{+}}\langle\hat{g}_{\omega}(t+\tau)\hat{v}_{\omega}(\tau),\hat{w}_{\omega}(t)\rangle\,\textrm{d}\tau\textrm{d}t$ (36)
@@
where $\hat{v}_{\omega},\hat{w}_{\omega}\in L_{2}(\mathbb{R}_{+},\mathbb{C}^{m})$ for almost all $\omega\in\mathbb{R}^{s}$. Therefore, $\mathcal{H}\geq 0$ because $\hat{\mathcal{H}}_{\omega}\geq 0$ for all $\omega\in\mathbb{R}^{s}$.
@@
due to (38) and the assumption that $\mathcal{H}$ is self-adjoint. The latter holds for all $\bar{v},\bar{w}\in L_{2}(\mathbb{R}_{+},\mathbb{C}^{m})$ and compact $\Omega\subset\mathbb{R}^{s}$, hence $\hat{\mathcal{H}}_{\omega}$ is self-adjoint for almost all $\omega\in\mathbb{R}^{s}$. Similarly, $\hat{\mathcal{H}}_{\omega}$ is nonnegative for almost all $\omega\in\mathbb{R}^{s}$. Due to [3, Theorem 5], this implies that $\hat{\Sigma}_{\omega}$ is of relaxation type and, thus, $\hat{g}_{\omega}$ is completely monotone for almost all $\omega\in\mathbb{R}^{s}$. Note that $\hat{g}_{\omega}(t)$ and its time derivatives are continuous in $\omega\in\mathbb{R}^{s}$ for each $t\in\mathbb{R}_{+}$ because the symbols $\hat{A}$, $\hat{B}$ and $\hat{C}$ are assumed to be continuous. Therefore, by continuity, $\hat{g}_{\omega}$ is completely monotone for all $\omega\in\mathbb{R}^{s}$, hence $\Sigma$ is of relaxation type due to Theorem 1. ∎